what equation describes the line that contains the points -2,5 and 3,2

Answer by stanbon(75887)

what equation describes the line that
contains the points -2,5 and 3,2
-----
slope = (2-5)/(3--2) = -3/5
------
intercept:
2 = (-3/5)(3) + b
---
2 = -9/5 + b
b = 19/5
------
y = (-3/5)x-(19/5)
-----------------------
Cheers,
Stan H.